OpenStudy (anonymous):

find the area of the rectangular heating-duct grill having a length of 2yds and a width of 21/4 yd. (type a whole number or mixed number)

OpenStudy (tkhunny):

Is the latter number \(\dfrac{21}{4}\;or\;2\dfrac{1}{4}\)?

OpenStudy (anonymous):

im not sure

OpenStudy (tkhunny):

Well, that's not very helpful. I would guess the latter is intended, but it's just not clear. Do both problems and let's see what you get. \(2\;yds\cdot \dfrac{21}{4}\;yds = ??\;yds^{2}\) \(2\;yds\cdot 2\dfrac{1}{4}\;yds = ??\;yds^{2}\)

OpenStudy (anonymous):

ok so how would i solve this

OpenStudy (tkhunny):

Multiply?